Calculate the ph of 100ml of 0.1 M HCl+ 100ml of 0.2 M H2So4+ 100 ml of 0.1M HNo3 +700ml of water?
 

Grade:11

1 Answers

Aashna
6 years ago
For calculating ph of mixture m1v1nf + m2v2nf2 + m3v3nf3 + m4v4nf4.= M(final)V(total)10+40+10 =M*1000 ( [H+] can be neglected as it is very very small compared to others. But volume will still be counted.)M = 6/100[H+] = .06-log[H+] = pH2 - log 61.23 =pH